Question 1: Consider a $4,700 deposit earning 9 percent interest per year for 9 years.

Part 1: What is the future value? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Part 2:  How much total interest is earned on the original deposit? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Part 3: How much is interest earned on interest? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Question 2.To borrow $1,400, you are offered an add-on interest loan at 7 percent. Three loan payments are to be made, one at four months, another at eight months, and the last one at the end of the year.

Compute the three equal payments. (Round your answer to 2 decimal places.)
Answer #1

Part 1: What is the future value

=4700*(1+9%)^9

=10207.90

Part 2:  How much total interest is earned on the original deposit

=10207.90-4700

=5507.90

Part 3: How much is interest earned on interest

=5507.90-(4700*9%*9)

=1700.90

the above is answer..